# Residual Diffusion Modeling for Km-scale Atmospheric Downscaling

### v3

@@ Abstract
Predictions of weather hazard require expensive km-scale simulations driven by coarser global inputs. Here, a cost-effective stochastic downscaling model is trained from a high-resolution 2-km weather model over Taiwan conditioned on 25-km ERA5 reanalysis. To address the multi-scale machine learning challenges of weather data, we employ a two-step approach Corrector Diffusion (CorrDiff), where a UNet prediction of the mean is corrected by a diffusion step. Akin to Reynolds decomposition in fluid dynamics, this isolates generative learning to the stochastic scales. CorrDiff exhibits skillful RMSE and CRPS and faithfully recovers spectra and distributions even for extremes. Case studies of coherent weather phenomena reveal appropriate multivariate relationships reminiscent of learnt physics: the collocation of intense rainfall and sharp gradients in fronts and extreme winds and rainfall bands near the eyewall of typhoons. Downscaling global forecasts successfully retains many of these benefits, foreshadowing the potential of end-to-end, global-to- km-scales machine learning weather predictions.
@@ 1 Introduction
Coarse-resolution 25-km global weather prediction is undergoing a machine learning renaissance thanks to autoregressive machine learning models trained on global reanalysis [6, 39, 11, 7, 30, 15, 14]. However, many applications of weather and climate data require kilometer-scale forecasts: e.g., risk assessment and capturing local effects of topography and human land use  [22]. Globally, applying ML at km-scale resolution poses significant challenges since training costs are superlinear with respect to the resolution of training data. Moreover, predictions from global km-scale physical simulators are not yet well tuned, so available training data can have worse systematic biases than coarse-resolution or established regional simulations [56, 27], and current data tends to cover short periods of time. Such datasets are also massive, difficult to transfer between data centers and frequently not produced on machines attached to significant AI computing resources like GPUs.
@@
In contrast, for regional simulation, scaling ML to km-scales is attractive. High-quality training data is abundant as many national weather agencies couple km-scale numerical weather models in a limited domain to coarser resolution global models [17] – a process called dynamical downscaling. Since these predictions are augmented by data assimilation from ground-based precipitation radar and other sensors, good estimates of regional km-scale atmospheric states exists [13]. Such dynamical downscaling is computationally expensive, which limits the number of ensemble members used to quantify uncertainties [38]. A common inexpensive alternative is to learn a statistical downscaling from these dynamical downscaling simulations and observations [60].
@@
In this context, ML downscaling enters as an advanced (non linear) form of statistical downscaling with potential to emulate the fidelity of dynamical downscaling. Several ML methods have been previously used for downscaling [9, 50, 19, 57, 38, 59, 2]. Convolutional Neural Networks, which reduce the input dimensions, have shown promise in globally downscaling climate (100km) data to weather scales (25km) [35, 47, 3, 45]. However, such deterministic ML approaches require interventions to produce useful probabilistic results, such as ensemble inference [47] or predicting the parameters of an assumed distribution [3]).
@@
The stochastic nature of atmospheric physics at km-scale [51] makes the downscaling inherently probabilistic, making it natural to apply generative models at these scales. Generative Adversarial Networks (GANs) have been tested, including for forecasting precipitation at km-scale in various regions [32, 43, 23, 46, 21, 59]; see the latter for a good review. Training GANs, however, poses several practical challenges including mode collapse, training instabilities, and difficulties in capturing long tails of distributions [61, 29, 48].
@@
Alternatively, diffusion models offer sample diversity and training stability [26, 16] alongside demonstrable skill in probabilistically generating km-scales. [1] used a diffusion model for predicting rain density in the UK from vorticity as an input, thus demonstrating potential for channel synthesis. [24] used a diffusion model for downscaling solar irradiance in Hawaii with a 1 day lead time, demonstrating the ability to simultaneously forecast. Moreover, diffusion models have been used directly for probabilistic weather forecasting and nowcasting [31, 33, 36]. See table S1 in 1 for more details.
@@
Building from this work, we turn to our challenge of interest – stochastically downscaling multiple variables simultaneously while also transferring input information to predict a new field (i.e. channel synthesis). If successful, this paves the way towards end-to-end ML downscaling systems that predicts regional high-resolution weather as a postprocessing of global predictions. As a proof of concept we will demonstrate such a ML model trained for the region surrounding Taiwan and show its downscaling abilities from both a global reanalysis and a global forecast model.
@@
    A physics-inspired, two-step approach (CorrDiff) to learn mappings between low- and high-resolution weather data with high fidelity.
@@
    CorrDiff provides realistic measures of stochasticity and uncertainty, in terms of Continuous Rank Probability Score (CRPS) and by comparing spectra and distributions.
@@
    CorrDiff reproduces the physics of coherent weather phenomena remarkably well, correcting frontal systems and typhoons.
@@
    CorrDiff is sample-efficient, learning effectively from just 4 years of data.
  5. 5.
    CorrDiff on a single GPU is at least 22 times faster and 1,300 times more energy efficient than the numerical model used to produce its high resolution training data, which is run on 928 CPU cores.
@@ 2 Generative downscaling: Corrector diffusion model
Consider a specific region on Earth, mapped onto a two-dimensional grid. Our input ${\bf y}\in\mathbbm{R}^{c_{\mathrm{in}}\times m\times n}$ is a low-resolution meteorological forecast taken from a 25-km global weather forecasting model (e.g., FourCastNet [39, 30, 7], or the Global Forecast System (GFS) [37]) . Here, $c_{\mathrm{in}}$ represents the number of input channels and $m,n$ represent the dimensions of a 2D subset of the globe. Our targets ${\bf x}\in\mathbbm{R}^{c_{\mathrm{out}}\times p\times q}$ come from corresponding data aligned in time $c_{\mathrm{out}}$ but having higher resolution, i.e. $p>m$ and $q>n$.
@@
In our proof of concept we use the ERA5 reanalysis as input, over a subregion surrounding Taiwan, with $m=n=36$, $c_{in}=20$ and $c_{out}=4$. See Table S2 for details about the inputs and outputs. The target data are 12.5 times higher resolution ($p=q=448$) and were produced using a radar-assimilating Weather Research and Forecasting (WRF) physical simulator [41] provided by the Central Weather Administration of Taiwan (CWA) [13] (i.e. CWA-WRF). Though imperfect, WRF is a SOTA model for km-scale weather simulations and is used operationally by several national weather agencies.
@@
The goal of probabilistic downscaling is to mimic the conditional probability density $p({\bf x}|{\bf y})$. To learn $p({\bf x}|{\bf y})$ we employ a diffusion model. Such models learn generative SDEs through the concept of score matching [26, 53, 28, 52, 5], with a forward and a backward processes working in tandem. In the forward process, noise is gradually added to the target data until the signal becomes indistinguishable from noise. The backward process then involves denoising the samples using a dedicated neural network to eliminate the noise. Through this sequential denoising process, the model iteratively refines the samples, bringing them closer to the target data distribution. The denoising neural network plays a critical role in this convergence, providing the necessary guidance to steer the samples towards accurate representations of the original data.
The motivation for CorrDiff is that naively applying conditional diffusion models to directly learn $p({\bf x}|{\bf y})$ was unsuccessful. The large uncertainty of variables such as radar reflectivity at these scales requires substantial noise levels during the forward process demands a large number of steps in backward process. This impedes learning and leads to poor sample fidelity [54].
@@
Figure 1: The workflow for training and sampling CorrDiff for generative downscaling. Top: Coarse-resolution global weather data at 25 km scale is used to first predict the mean $\boldsymbol{\mu}$ using a regression model, which is then stochastically corrected using EDM diffusion ${\bf r}$, together producing the probabilistic high-resolution 2 km-scale regional forecast. Bottom right: diffusion model is conditioned with the coarse-resolution input to generate the residual ${\bf r}$ after a few denoising steps. Bottom left: the score function for diffusion is learned based on the UNet architecture.
@@
where ${\bf y}$ and ${\bf x}$ are the input and target respectively. This signal decomposition is inspired by Reynolds decomposition in fluid-dynamics [40] and climate data analytics. Assuming the regression learns the conditional mean accurately, i.e., ${\bf\mu}\approx\mathbbm{E}[{\bf x}|{\bf y}]$, the residual is zero mean, namely $\mathbbm{E}[{\bf r}|{\bf y}]\approx 0$, and as a result ${\rm var}({\bf r}|{\bf y})={\rm var}({\bf x}|{\bf y})$. Accordingly, based on law of total variance [10], one can decompose the variance as
@@
That is, the residual formulation ensures targets that have significantly smaller variance. According to (2), the variance reduction is more pronounced when ${\rm var}(\mathbbm{E}[{\bf x}|{\bf y}])$ is large, e.g., in the case of Typhoons. For our specific target data we find that the actual variance reduction is significant, especially at large scales; see section 3 and Figure S2. Moreover, the difference $\bf r$ is very localized and has no significant spatial auto-correlation at ranges beyond 30 km. In contrast, $\bf x$ has significant auto-correlation even at a 200 km distance. To sum it up, the main idea of CorrDiff is that learning the distribution $p(\bf r)$ can be much easier than learning the distribution $p(\bf x)$. Since modeling multi-scale interactions is a daunting task in many physics domains, we expect this approach could be widely applied. More details are described in Section 5 and the outline is depicted in Fig. 1.
Figure 2: Power spectra and distributions for the interpolated ERA5 input, CorrDiff, RF, Reg, and WRF. These results reflect reductions over space, time and for CorrDiff across 192 different samples. Left: Power spectra for kinetic energy (top), 2 meter temperature (middle) and radar reflectivity (bottom). Right: distributions of windspeed, (top), 2 meter temperature (middle) and radar reflectivity (bottom). Radar reflectivity is not included in the ERA5 dataset. We show the log-PDF to highlight the differences at the tails of the distributions. Here wavenumber is the inverse of a wavelength.
@@ 3 Results
The target data (WRF) span four years (2018-2021) and additional first four months of 2022, at one hour time resolution. We use the first 3 years for training and the rest for testing. The input (course resolution) data is taken from the ERA5 reanalysis for the corresponding dates. See 5 and S2 for details.
CorrDiff downscaling is compared with the input and target data and several baseline models, for deterministic and probabilistic skill scores, distributions and power spectra, and case studies of various coherent atmospheric phenomena.
@@ 3.1 Skill
Compared with several baselines, CorrDiff shows the highest deterministic skill (MAE) followed by the UNet regression (Reg), the random forest (RF) and the interpolation of ERA5 (1). The difference between MAE of CorrDiff and that of the UNet reflects the correction of the sample mean by the diffusion step. The consistent improvement in MAE for all the target variables between the UNet and CorrDiff suggests that the generative CorrDiff model (step 2) can correct some of the remaining biases after the UNet-regression (step 1).
Method Metric u10m Radar v10m t2m
CorrDiff CRPS 0.29 0.53 0.31 0.14
CorrDiff MAE 0.40 0.74 0.43 0.19
UNet-regression MAE 0.45 0.77 0.50 0.24
Random forest regression MAE 1.15 3.58 1.28 0.81
ERA5 bilinear interpolation MAE 1.18 - 1.28 0.97
Table 1: Skill scores evaluated from 204 date and time combinations taken randomly from the out-of-sample year (2021). For CorrDiff each evaluation used 192 samples. The table compares CorrDiff, UNet-regression, random forest, and interpolated ERA5 predictions in terms of MAE of the ensemble mean and CRPS for different atmospheric variables. For deterministic predictions, MAE and CRPS are equivalent. A separate RF is fit with scikit-learn for each of the 4 output channels with 100 trees and the default hyperparameters. While crude, this RF provides a simple (and easily tuned) baseline for the performance of Reg.
@@ 3.2 Spectra and distributions
CorrDiff is able to predict the correct power spectra of kinetic energy (KE), 2-meter temperature and radar reflectivity compared with the baselines in Table 1 (Fig. 2 a-c). The benefits of CorrDiff relative to the UNet baseline are stronger for some predicted variables than others. Improvements are most striking for radar reflectivity (Fig. 2 c) for which CorrDiff matches the target power spectra while the UNet-regression is significantly worse. This is expected as radar reflectivity comes from precipitation, which is linked to intrinsically stochastic physics (see [59]), i.e. requires the generation step for skillful prediction. In contrast, for 2-meter temperature (Fig. 2 b) the performance of UNet-regression is comparable to CorrDiff; temperature downscaling is expected to be mostly driven by sub-grid variations in topography that can be learned determinstically from the static grid embeddings.
Good skill is also found when looking at the corresponding probability distributions (Fig. 2 d-f). CorrDiff is able to match the target distribution, including the heavy-tailed structure, which is significantly improved from the the prediction based on the UNet alone. These rare extreme values reflect most of the risk. For windspeed (Fig. 2 d), the UNet and the baseline models underestimate the probability of winds faster than 20 m/s. For 2-meter temperatures (Fig. 2 e), the baseline models overestimate the probability of warm extremes and underestimate the probability of cold extremes. Radar reflectivity proves to be the most challenging distribution to reproduce. CorrDiff outperforms the other models in reproducing the reflectivity PDF (Fig. 2f). Although it overestimates the occurrence of reflectivity at lower values (i.e. weakly precipitating systems), the match of slopes at high values attests that the power-law relationship regulating the heavy-tail structure of rain extremes is emulated skillfully.

### v4

# Residual Corrective Diffusion Modeling for Km-scale Atmospheric Downscaling

@@ Abstract
The state of the art for physical hazard prediction from weather and climate requires expensive km-scale numerical simulations driven by coarser resolution global inputs. Here, a generative diffusion architecture is explored for downscaling such global inputs to km-scale, as a cost-effective machine learning alternative. The model is trained to predict 2km data from a regional weather model over Taiwan, conditioned on a 25km global reanalysis. To address the large resolution ratio, different physics involved at different scales and prediction of channels beyond those in the input data, we employ a two-step approach where a UNet predicts the mean and a corrector diffusion (CorrDiff) model predicts the residual. CorrDiff exhibits encouraging skill in bulk MAE and CRPS scores. The predicted spectra and distributions from CorrDiff faithfully recover important power law relationships in the target data. Case studies of coherent weather phenomena show that CorrDiff can help sharpen wind and temperature gradients that co-locate with intense rainfall in cold front, and can help intensify typhoons and synthesize rain band structures. Calibration of model uncertainty remains challenging. The prospect of unifying methods like CorrDiff with coarser resolution global weather models implies a potential for global-to-regional multi-scale machine learning simulation.
@@ 1 Introduction
Coarse-resolution 25-km global weather prediction is undergoing a machine learning renaissance with the recent advance of autoregressive machine learning models trained on global reanalysis [6, 46, 12, 7, 36, 17, 16, 51, 11, 37]. However, many applications of weather and climate data require kilometer-scale forecasts: e.g., risk assessment and capturing local effects of topography and human land use  [25]. Globally, applying ML at km-scale resolution poses significant challenges since training costs are superlinear with respect to the resolution of training data. Moreover, predictions from global km-scale physical simulators are not yet well tuned, so available training data can have worse systematic biases than coarse-resolution or established regional simulations [66, 31], and current data tends to cover short periods of time. Such datasets are also massive, difficult to transfer between data centers and frequently not produced on machines attached to significant AI computing resources like GPUs.
@@
In contrast, for regional simulation, using ML to conditionally generate km-scales is attractive. High-quality training data are available as many national weather agencies couple km-scale numerical weather models in a limited domain to coarser resolution global models [20] – a process called dynamical downscaling. Since these predictions are augmented by data assimilation from ground-based precipitation radar and other sensors, good estimate of regional km-scale atmospheric states exists [15]. Such dynamical downscaling is computationally expensive, which limits the number of ensemble members used to quantify uncertainties [45].
A common inexpensive alternative is to learn a statistical downscaling from these dynamical downscaling simulations and observations [72]. This is typically done by learning the values of several parameters of a statistical mapping (e.g. quantile mapping, generalized linear regression) that best match a regional high resolution data [3]. In this context, ML downscaling enters as an advanced (non linear) form of statistical downscaling [54] with potential to exceed the fidelity of conventional statistical downscaling.
@@
Several ML methods have been previously used for downscaling [9, 60, 22, 68, 45, 71, 2]. Convolutional Neural Networks have shown promise in globally downscaling climate (100km) data to weather scales (25km) [42, 56, 3, 53]. However, such deterministic ML approaches require interventions to produce useful probabilistic results, such as ensemble inference [56] or predicting the parameters of an assumed distribution [3]).
@@
The stochastic nature of atmospheric physics at km-scale [61] renders downscaling inherently probabilistic, making it natural to explore generative models at these scales. Generative Adversarial Networks (GANs) have been tested, including for forecasting precipitation at km-scale in various regions [39, 50, 26, 55, 24, 71]; see the latter for a good review. Training GANs, however, poses several practical challenges including mode collapse, training instabilities, and difficulties in capturing long tails of distributions [73, 35, 58].
@@
Alternatively, diffusion models offer training stability [29, 18] alongside demonstrable skill in probabilistically generating km-scales. [1] used a diffusion model for predicting rain density in the UK from vorticity as an input, thus demonstrating potential for channel synthesis. [27] used a diffusion model for downscaling solar irradiance in Hawaii with a 1 day lead time, demonstrating the ability to simultaneously forecast. Moreover, diffusion models have been used directly for probabilistic weather forecasting and nowcasting [38, 40, 43, 67] – including global ensemble predictions that outperform conventional weather prediction on a range of important stochastic metrics at 0.25-degree resolution [51]. See table S1 in 1 for more details.
@@
Building upon these works, we turn to our challenge of interest – stochastically downscaling multiple variables simultaneously while also transferring input information to predict a new field (i.e., channel synthesis). If successful, this paves the way towards ML downscaling systems that produce regional high-resolution weather as a postprocessing of coarser global predictions. As a proof of concept we will demonstrate such a ML model trained for the region surrounding Taiwan.
@@
    A physics-inspired, two-step approach (CorrDiff) to simultaneously learn mappings between low- and high-resolution weather data across multiple variables with high fidelity alongside new channel synthesis.
@@
    For the case studies considered, CorrDiff adds physically realistic improvements to the representation of under-resolved coherent weather phenomena – frontal systems and typhoons.
@@
    CorrDiff is sample-efficient, learning effectively from just 3 years of data.
@@
    CorrDiff on a single GPU is at least 22 times faster and 1,300 times more energy efficient than the numerical model used to produce its high-resolution training data, which is run on 928 CPU cores, see 6.3 for details.
@@ 2 Generative downscaling: Corrector diffusion model
Consider a specific region on Earth, mapped onto a two-dimensional grid. Our input ${\bf y}\in\mathbbm{R}^{c_{\mathrm{in}}\times m\times n}$ is a low-resolution meteorological data taken from a 25-km global reanalysis, or weather forecasting model (e.g., FourCastNet [46, 36, 7], or the Global Forecast System (GFS) [44]). Here, $c_{\mathrm{in}}$ represents the number of input channels and $m,n$ represent the dimensions of a 2D subset of the globe. Our targets ${\bf x}\in\mathbbm{R}^{c_{\mathrm{out}}\times p\times q}$ come from corresponding data aligned in time $c_{\mathrm{out}}$ but having higher resolution, i.e., $p>m$ and $q>n$.
@@
In our proof of concept we use the ERA5 reanalysis as input, over a subregion surrounding Taiwan, with $m=n=36$, $c_{in}=12$ and $c_{out}=4$. See Table S2 for details about the inputs and outputs. The target data are 12.5 times higher resolution ($p=q=448$) and were produced using a radar-assimilating Weather Research and Forecasting (WRF) physical simulator [48] provided by the Central Weather Administration of Taiwan (CWA) [15] (i.e., CWA-WRF), which employs a dynamical downscaling approach. Though imperfect, WRF is a SOTA model for km-scale weather simulations and is used operationally by several national weather agencies.
@@
The goal of probabilistic downscaling is to mimic the conditional probability density $p({\bf x}|{\bf y})$. To learn $p({\bf x}|{\bf y})$ we employ a diffusion model. Such models learn stochastic differential equations (SDEs hereafter) through the concept of score matching [29, 63, 33, 62, 5], with a forward and a backward processes working in tandem. In the forward process, noise is gradually added to the target data until the signal becomes indistinguishable from noise.
The backward process then involves denoising the samples using a dedicated neural network to eliminate the noise. Through this sequential denoising process, the model iteratively refines the samples, bringing them closer to the target data distribution. The denoising neural network plays a critical role in this convergence, providing the necessary guidance to steer the samples towards accurate representations of the original data.
The development of CorrDiff was motivated by the limitations observed when using conditional diffusion models to directly learn $p(\mathbf{x}|\mathbf{y})$. This approach showed slow convergence and resulted in poor-quality images with incoherent structures. This was surprising because conditional diffusion models have been successfully applied to super-resolution tasks in natural image restoration, as demonstrated in works like [57]. We hypothesize that the significant distribution shift between the input variables and challenging target variables, particularly the 1-hour maximum derived radar reflectivity (hereafter referred to as radar reflectivity), necessitates high noise levels during the forward process and numerous steps in the backward process. Our experiments indicated that these requirements hindered learning and compromised sample fidelity [64]. This issue is particularly relevant for the downscaling task, which must account for large spatial shifts, correct biases in static features like topography, and synthesize entirely new channels like radar reflectivity. By comparison, the task of super-resolution in natural images is much simpler, as it typically involves local interpolation and does not face the same level of distributional challenges.
@@
Figure 1: The workflow for training and sampling CorrDiff for generative downscaling. Top: Coarse-resolution global weather data at 25 km scale is used to first predict the mean $\boldsymbol{\mu}$ using a regression model, which is then stochastically corrected using an Elucidated Diffusion Model (EDM) ${\bf r}$, together producing the probabilistic high-resolution 2 km-scale regional forecast. Bottom right: diffusion model is conditioned with the coarse-resolution input to generate the residual ${\bf r}$ after a few denoising steps. Bottom left: the score function for diffusion is learned based on the UNet architecture.
@@
where ${\bf y}$ and ${\bf x}$ are the input and target respectively. This signal decomposition is inspired by Reynolds decomposition in fluid-dynamics [47] and climate data analytics. Assuming the regression learns the conditional mean accurately, i.e., ${\bf\mu}\approx\mathbbm{E}[{\bf x}|{\bf y}]$, the residual is zero mean, namely $\mathbbm{E}[{\bf r}|{\bf y}]\approx 0$, and as a result ${\rm var}({\bf r}|{\bf y})={\rm var}({\bf x}|{\bf y})$. Accordingly, based on the law of total variance [10], one can decompose the variance as
@@
That is, the residual formulation reduces the variance of the target distribution. According to (2), the variance reduction is more pronounced when ${\rm var}(\mathbbm{E}[{\bf x}|{\bf y}])$ is large, e.g., in the case of typhoons. For our specific target data we find that the actual variance reduction is significant, especially at large scales; see section 3 and Figure S3. To sum it up, the main idea of CorrDiff is that learning the distribution $p(\bf r)$ can be much easier than learning the distribution $p(\bf x)$. Since modeling multi-scale interactions is a daunting task in many physics domains, we expect this approach could be widely applied. More details are described in Section 5 and the outline is depicted in Fig. 1.
Figure 2: Power spectra and distributions for the interpolated ERA5 input, CorrDiff, RF, UNet, and WRF. These results reflect reductions over space, time and for CorrDiff across 32 different samples per each time. Left: Power spectra for kinetic energy (top), 2-meter temperature (middle) and radar reflectivity (bottom). Right: distributions of windspeed, (top), 2-meter temperature (middle) and radar reflectivity (bottom). Radar reflectivity is not included in the ERA5 dataset. We show the log-PDF to highlight the differences at the tails of the distributions. Here wavenumber is the inverse of a wavelength.
Our target (WRF) dataset spans 2018 through 2021 at hourly time resolution. We use 2018 through 2020 for training and the rest for testing. We additionally use several days of typhoon data from 2023 and some snapshots of a coherent frontal weather system from 2022 for testing case studies. The input (coarse resolution) data are taken from the ERA5 reanalysis for the corresponding times. The UNet and a random forest are used as baselines. See Section 5 and Table S2 for details.
@@ 3 Results
In this section CorrDiff downscaling is compared with the input and target data as well as with several baseline models. A common set of 205 randomly selected out-of-sample date and time combinations from 2021 is used for computing metrics and spectra and for intercomparing CorrDiff with the baseline models. For CorrDiff ensemble predictions are examined using a 32-member ensemble; larger ensembles do not meaningfully modify the key findings below (not shown).
3.1 Baseline Models
As baselines, we use an interpolation of the condition data (ERA5), a Random Forest (RF) and the regression step of CorrDiff (UNet). Using the same 12 low-resolution input channels we fit an RF for each of the 4 output channels with 100 trees and the default hyperparameters. The RF is applied independently at each horizontal location similar to a $1\times 1$ convolution. While crude, this RF provides a simple (and easily tuned) baseline for the performance of the UNet. To ensure the best performance for each channel individually, we train separate RFs for each output channel.
@@ 3.2 Skill
When comparing the CRPS of CorrDiff with the MAE of the UNet and the other baselines, CorrDiff exhibits the most skill, followed by the UNet, the random forest (RF) and the interpolation of ERA5 (Table (1). The slight degradation in MAE of CorrDiff compared to that of the UNet is expected, as the diffusion model optimizes the Kullback–Leibler divergence as opposed to optimizing for MAE loss optimized by the UNet (see Section 5.2.2). In table S3 we show that pooled scores (CRPS and MAE) tell a similar story.
Radar t2m u10m v10m
CorrDiff (CRPS) 1.90 0.55 0.86 0.95
CorrDiff (MAE) 2.54 0.65 1.08 1.19
UNet 2.51 0.64 1.10 1.21
RF 3.56 0.81 1.14 1.26
ERA5 - 0.97 1.17 1.27
Table 1: MAE and CRPS scores evaluated from 205 date and time combinations taken randomly from the out-of-sample year (2021). For CorrDiff the CRPS was computed using 32 ensemble members and the MAE is computed for the ensemble mean. For deterministic predictions given by all other models, MAE and CRPS are equivalent. The differences between CorrDiff, UNet, and RF are all statistically significant (see SI Section 6.4). CorrDiff has lower CRPS than the UNet in 205/205 of the validation times.
@@ 3.3 Spectra and distributions
Relative to deterministic baselines, CorrDiff significantly improves the realism of power spectra for 10-meter kinetic energy (KE), 2-meter temperature and synthesized radar reflectivity. Variance missing from the UNet is restored by the corrective diffusion (blue-dashed vs blue-solid) – especially for the radar reflectivity channel at all length scales (Fig. 2c), but also for kinetic energy between 10–200 km length scales (Fig. 2a) and to a lesser extent for temperature on 10–50 km length scales. Temperature downscaling is an easier task that is expected to be mostly driven by sub-grid variations in topography that can be learned deterministically from the static grid embeddings. Evidently, synthesizing radar from only indirectly related inputs is the task that most benefits from the corrective diffusion component of CorrDiff.
This is corroborated by analysis of probability distributions (Fig. 2d-f) – for the radar reflectivity channel both the UNet and RF fail to produce realistic statistics, but CorrDiff is able to match the target distribution between 0 and 43 dbz while significantly improving on the UNet (Fig. 2f). In contrast to the radar channel, the hot and cold tails of the CorrDiff-generated surface temperature distribution are only incrementally improved relative to the UNet (Fig. 2e) and the overall windspeed PDF is virtually unchanged relative to the UNet, despite the scale-selective variance enhancements noted previously. Overall, CorrDiff produces encouraging probability distributions, with the caveat that apparent agreement of generated tail structures should be viewed as provisional given that our chosen validation sample of 205 independent calendar times imperfectly samples especially low likelihood/high-impact extremes.
While encouraging, CorrDiff’s emulation of radar statistics is also imperfect; generated radar variance is somewhat under-estimated on length scales greater than 100 km and over-estimated for the 10–50 km length scales (Fig. 2c), associated with an overall overdispersive PDF (Fig. 2f).
3.4 Model Calibration
Analysis of the ensemble spread of our 32-member CorrDiff predictions shows they are not yet optimally calibrated. Figure 3 demonstrates that the predictions are overall under-dispersive for most channels – the ensemble spread is too small relative to ensemble mean error and rank histograms indicate that observed values frequently fall above or below the range of predicted values. Optimizing the stochastic calibration of CorrDiff is a logical area for future development.
Figure 3: Evaluation of model calibration base do the same validation set used in figure 2 and 1. Left column - the ensemble standard deviation as a function of the RSME of mean prediction for the 4 channels. The standard deviation is adjusted with a factor $\sqrt{(}1+1/n)$ so that a ratio of one represents a perfectly tuned model. Right column shows the corresponding rank histograms.